You own a portfolio that has $1,950 invested in Stock A and $3,250 invested in Stock B. If the expected returns on these stocks are 8 percent and 16 percent, respectively, what is the expected return on the portfolio?

Multiple Choice 12.00% 13.00% 11.00% 13.26% 13.65%
Answer #1

Total investment = 1950 + 3250 = 5200

Weight of stock A = 1950/5200 = 0.375

Weight of stock B = 1 - 0.375 = 0.625

The expected return = 0.375 * 0.08 + 0.625 * 0.16

The expected return = 0.13 = 13%

Can you please upvote? Thank You :-)